Find the eccentric angles of the extrimities of latus recta of the ellipse x^2/a^2+y^2/b^2=1(a>b)

Grade:11

2 Answers

Piyush Kumar Behera
7 years ago
 
The x coordinate of latus rectum is the x coordinate of the  ends of the latus rectum =\pmae
And the parametric coordinates of ellipse are  (acos\theta,bsin\theta)
 
Now,if the eccentric angle of ends of latus rectum is \theta then acos\theta=+ae
\theta=cos-1e
Manish
7 years ago
use the formula
this is the most imp thing you need to know...............
x=acos\theta y=bsin\theta \frac{y}{x}=\frac{b}{a}tan\theta
tan\theta=ay/bx
where ends of latus recta are (\pmae,b2/a)
i.e x=ae, y=b2/a
\thereforetan\theta =(a*b^{2}/a)/(b*ae) =b/ae